The numerator of a certain fraction is 3 less than the denominator. If the numerator is tripled and the denominator is increased by 7, the resulting fraction is equal to 3/2. What was the original fraction?
Answer by richwmiller(17219)

3*(x-3)/(x+7)=3/2
x=13
30/20
o=(x-3)/x
o=10/13
original fraction =10/13